HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

INFLUENCE OF THE COLLISION ENERGY ON THE REACTION RATE OF THE D_{2}^{+} + H_{2}O ION-MOLECULE REACTION NEAR 0 K

\underline{D. SCHLANDER} , R. HAHN, V. ZHELYAZKOVA, F. MERKT, Institute for Molecular Physical Science, ETH Zürich, Vladimir-Prelog-Weg 2, 8093 Zürich,

Switzerland

Ion-molecule reactions are important reactions in atmospheric chemistry, astrophysics, and plasma physics.^{1} Reliable values for the reaction rates are key to modelling the relevant reaction networks. This poster focusses on the reactions of water with D_{2}^{+} at the low temperatures in the range of 0-50 K of interstellar molecular clouds.^{1}^{,}^{2} For technical reasons D_{2}^{+} is used instead of H_{2}^{+}.

Because ions are easily accelerated by (stray) electric fields, high Rydberg states (n ≈ 30, v^{+} = 0 , N^{+} = 0) are utilised as proxies for the ionic reaction partner, because the Rydberg-electron does not significantly affect the reaction but shields it from electric fields.^{3}^{,}^{4} We use a merged-beam approach featuring a Rydberg-Stark deflector and decelerator to access the collision-energy range below k_{B} · 10 K. The molecular beam of water is characterised.

The reaction rates are modelled with an adiabatic capture model inspired by earlier work in Refs. ^{5}^{,}^{6}^{,}^{7} which is constructed by modifying the Langevin model with the rotational-state-specific energy of the neutral molecules in the field of the ion.^{5}^{,}^{6}^{,}^{7}^{,}^{8} For Maxwell-Boltzmann rotational-state distributions below 100 K the model predicts an increase of the reaction-rate constant at low collision energies.

1doi:10.1038/s42254-020-0198-0, B. A. McGuire, O. Asvany, S. Brünken, S. Schlemmer, Nat. Rev.

Phys., 2, 402-410, (2020).

2doi:10.1146/annurev.pc.46.100195.000331, E. Herbst, Annu. Rev. Phys. Chem., 46, 27-53, (1995).

3doi:10.1103/PhysRevA.90.043403, P. Allmendinger, J. Deiglmayr, J. A. Agner, H. Schmutz, F.

Merkt, Phys. Rev. A, 90, 043403, (2014).

4doi:10.1088/1367-2630/ac231d, F. B. V. Martins, V. Zhelyazkova, C. Seiler, F. Merkt, New J. Phys., 23, 095011, (2021).

5doi:10.1063/1.453701, J. Troe, Chem. Phys., 87, 2773-2780, (1987).

6doi:10.1080/00268978500100461, D. C. Clary, Mol. Phys., 54, 605-618, (1985).

7doi:10.1063/1.472479, J. Troe, J. Chem. Phys., 105, 6249-6262, (1996).

8doi:10.1039/D1CP03116C, V. Zhelyazkova, F. B. V. Martins, J. A. Agner, H. Schmutz, F. Merkt, Phys. Chem. Chem. Phys.

p-number: p060  Submitted on Fri, 09 Jun 2023 16:03:58 +0200